Solutions of:
2X^2-5X-3=0

Answer by MathmaticsMaster(2)

2X^2 - 5X - 3 = 0 Original problem.
(2X + 1)(X - 3) = 0 Factor it completely.
2X + 1 = 0 X - 3 = 0 Separate it into all of the parts that equal zero.
2X = -1 X = 3 Finish.
X = - 1/2 X = 3 Finish.
Your answers are -1/2 and 3.